### v1

@@ Abstract
Accurate yet transferable machine-learning interatomic potentials (MLIPs) are essential for accelerating materials and chemical discovery. However, most universal MLIPs overfit to narrow datasets or computational protocols, limiting their reliability across chemical and functional domains. We introduce a transferable multi-domain training strategy that jointly optimizes universal and task-specific parameters through selective regularization, coupled with a domain-bridging set (DBS) that aligns potential-energy surfaces across datasets. Systematic ablation experiments show that small DBS fractions (0.1%) and targeted regularization synergistically enhance out-of-distribution generalization while preserving in-domain fidelity. Trained on fifteen open databases spanning molecules, crystals, and surfaces, our model, SevenNet-Omni, achieves state-of-the-art cross-domain accuracy, including adsorption-energy errors below 0.06 eV on metallic surfaces and 0.1 eV on metal-organic frameworks. Despite containing only 0.5% r2SCAN data, SevenNet-Omni reproduces high-fidelity r2SCAN energetics, demonstrating effective cross-functional transfer from large PBE datasets. This framework offers a scalable route toward universal, transferable MLIPs that bridge quantum-mechanical fidelities and chemical domains.
@@ 1 Introduction
By enabling large-scale atomistic simulations with accuracy approaching that of ab initio methods such as density functional theory (DFT), machine-learning interatomic potentials (MLIPs) have significantly expanded the scope of computational materials science [1, 2, 3, 4, 5]. Recently, pretrained universal MLIPs (uMLIPs) have attracted considerable attention because they can bypass the time-consuming creation of ab initio datasets for model training [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The large databases and deep neural-network architectures behind these models enable consistent accuracy outside the training domain [16, 17]. Currently, a multitude of large-scale ab initio databases are publicly available for training uMLIPs, covering most material classes [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Representative databases are shown in Fig. 1a, which indicates that the databases usually focus on specific chemical domains, reflecting distinct community interests. Each database is obtained with a unique set of computational protocols such as exchange-correlation (XC) functional, program, and ionic potentials.
@@
As materials engineering advances, the demand for realistic simulations that span multiple chemical domains is rapidly increasing. Examples include catalytic reactions on metal surfaces in aqueous environments [34, 35]; atomic layer deposition (ALD) in semiconductor processing, where metalorganic precursors decompose on metal or dielectric substrates [36]; the formation and evolution of the solid-electrolyte interphase [37, 38, 39]; and polymer-inorganic composites [40]. These simulations would require uMLIPs that can deliver consistent accuracy across distinct chemical domains by being trained over multiple databases. As mentioned earlier, each database is constructed with different computational protocols. Simply combining these databases, even with element-dependent energy shifts, would introduce significant noises. Therefore, it is necessary to employ multi-task training that separately respects the characteristic potential energy surface (PES) represented by each database. So far only a few uMLIPs such as DPA-3.1 [9] and UMA [8], are trained with the diverse databases concurrently within the multi-task framework.
@@
Because uMLIPs must reproduce energies consistent with a chosen set of computational parameters (e.g., XC functional or basis set), transfer learning to bridge differences across computational settings and chemical domains is essential for training uMLIPs for multi-domain applications. This challenge is particularly pronounced in simulations involving both inorganic compounds and small molecules, which are common in real-world applications. In addition to large structural differences, databases for inorganic materials are generated with semilocal functionals, whereas those for small molecules are typically generated with hybrid functionals (Fig. 1a). Since semilocal functionals are often preferred for combined systems [41, 42], transfer learning across chemical domains and XC functionals must be ensured. Most prior work, including widely used benchmarks, has evaluated multi-task uMLIPs within single-domain scenarios rather than in complex multi-domain applications [43, 9, 8]; consequently, efficient methods for maximizing cross-domain knowledge transfer remain underexplored.
@@ 2.1 Training strategies for cross-domain generalization
In this study, we address the training of MLIPs using multiple ab initio databases that differ both in the types of atomic configurations (e.g., inorganic versus organic systems) and in the underlying computational approaches (e.g., semilocal versus hybrid functionals). To construct a single MLIP capable of accommodating databases generated under heterogeneous computational settings, we employ a multi-task MLIP framework, where each task corresponds to a different database. (In our original paper, we referred to this approach as multi-fidelity [44]. However, since the present training datasets often share the same level of fidelity, such as Perdew-Burke-Ernzerhof (PBE) [45] and revised PBE (RPBE) [46], we now use the term multi-task for the same methodological approach.) Within this framework, the model parameters are divided into two categories: (i) shared parameters () that are universally applied across all databases, and (ii) task-specific parameters () that are optimized exclusively for the database corresponding to task . Formally,
@@
collects higher-order terms. Equation 2 thus separates the contributions that do and do not depend on . On the right-hand side, depends solely on the shared parameters and is therefore referred to as the common PES. The second term represents the task-specific contribution.
@@
To address this issue, we introduce regularization on . By penalizing large values of , regularization encourages the model to rely more heavily on the common PES, ensuring that the shared representation captures the essential bonding characteristics across both databases. This is illustrated schematically in the central panel of Fig. 1c, where the common PES now remains close to the ground-truth PES of both tasks, thereby enhancing knowledge transfer between the two databases.
While the regularization of is effective in improving transfer learning, it tends to fit the PES to the ground truth of in the domain. This limitation arises from the complete absence of at . To address this, we introduce a domain-bridging set (DBS), which augments by evaluating a subset of with the ab initio method used for (see the right panel of Fig. 1c). Incorporating a DBS requires only marginal computational effort (approximately 0.1% of the original database) yet yields substantial improvements in the accuracy of task at .
To demonstrate the effectiveness of the proposed approach, we employ the MPtrj [18] and SPICE [28] databases for MLIP training. The MPtrj database, representative of inorganic bulk systems, is generated using the PBE functional, whereas the SPICE database, curated for organic molecules, is based on hybrid functional of B97M [47]. For multi-task modeling, we adopt the SevenNet-MF architecture [44] with multi-fidelity capability and evaluate model performance on organic molecules at the PBE level by computing the force mean absolute error (MAE) relative to PBE reference results for the SPICE structures.
The evaluation results are presented in Fig. 1d. As a baseline, we trained a single-task model using only the MPtrj database (left region), which exhibits large errors for molecular structures evaluated at the PBE level (denoted as Molecule@PBE). We then examined the impact of different regularization strategies in SevenNet-MF training, comparing a multi-task MLIP without regularization, a model trained with conventional regularization applied to both and , and a model trained with regularization applied only to . As shown in the middle region of Fig. 1d, all models achieve comparable force accuracy within their respective training domains and computational settings (Crystal@PBE and Molecule@Hybrid), regardless of the regularization scheme. However, the force MAE differs significantly when evaluated on out-of-domain organic molecular configurations at the PBE level (Molecule@PBE), while all multi-task MLIPs outperform the single-task baseline. In particular, selective regularization of yields greater improvements than conventional regularization applied to all model parameters, underscoring its advantage for optimizing cross-database transferability. The right axis shows that model accuracy systematically improves as the L2 norm of decreases, indicating that effective knowledge transfer is facilitated by suppressing task-specific contributions.
@@
Building on these observations, we developed a uMLIP, 7net-Omni, by concurrently training the SevenNet-MF architecture on 15 publicly available databases (covering 13 different computational methods) with selective task regularization and DBS. The training set spans a diverse collection of ab initio databases encompassing molecular, crystalline, metal-organic framework (MOF), and surface systems, as schematically illustrated in Fig. 1a. The specific databases, along with their chemical domains and XC functionals, are listed in Table 1. The present uMLIP concerns the charge neural system so charged structures in SPICE, OMol25 and QCML are excluded. Each database is assigned to a distinct task, but databases with identical computational protocols are grouped under the same task. We also use the term channel to refer to task-specific inference. Further details of dataset composition, sampling ratios, and training procedures are provided in the Methods section. To construct DBS, we adopted the computational settings used in the MPtrj database and performed single-point calculations on approximately 0.1% of structures from six representative databases (see Methods for details of the sampling criterion). The numbers of sampled structures from each database are summarized in Table S1.
@@
To examine whether the model encodes computational methods in an explainable way, we analyzed the trained parameters of 7net-Omni associated with task embeddings. The latent vectors were visualized using principal component analysis (PCA) in Fig. 1b. We find that the task-dependent weights form distinct clusters that primarily correlate with the choice of XC functional and the application of a Hubbard correction. For example, molecular databases employing hybrid functionals group together in a similar region. This indicates that the model captures both similarities and discrepancies among different levels of theory, thereby contributing to the enhanced generalizability of the multi-task framework. Comparable analyses have been reported in the context of multi-task property prediction using the multiXC-QM9 dataset [48, 49]. However, unlike the present observation, it produced mixed results; for instance, embeddings generated with the same XC functional but different basis sets exhibited low similarity. This difference may stem from the use of selective task regularization in 7net-Omni, which allows the model to better recognize cross-theory similarities.
@@
| Database abbreviation | Task (Channel) | Number of structures | Domain | XC functional | Over- sampling | Reference |
| MPtrj | mpa | 1,580,394 | Inorganic crystal | PBE(+) | 15 | [18] |
| Alex1 | mpa | 12,069,066 | Inorganic crystal | PBE(+) | 4 | [19, 20, 21] |
| DBS | mpa | 121,141 | General | PBE(+) | 15 | - |
| OMat24 | omat24 | 101,901,967 | Inorganic crystal | PBE(+) | 1 | [22] |
| MatPES | matpes | 412,976 | Inorganic crystal | PBE | 40 | [23] |
| OC202 | oc20 | 30,757,554 | Catalyst (metal) | RPBE | 2 | [24] |
| OC22 | oc22 | 8,210,850 | Catalyst (oxide) | PBE(+) | 6 | [25] |
| ODAC23 | odac23 | 4,082,662 | MOF | PBE-D3 | 1 | [26] |
| OMol25 (low)3 | omol25 | 60,852,034 | Molecule | B97M-V | 1 (5)3 | [27] |
| OMol25 (high)3 | omol25_high | 1,390,125 | Molecule | B97M-V | 5 | [27] |
| SPICE4 | spice | 1,738,530 | Molecule | B97M | 5 | [28] |
| QCML | qcml | 18,301,834 | Molecule | PBE0 [50] | 1 | [29] |
| MAD | mad | 86,049 | General | PBEsol [51] | 100 | [30] |
| MP-r2SCAN | mp_r2scan | 50,231 | Inorganic crystal | r2SCAN | 40 | [52, 53] |
| MatPES-r2SCAN | matpes_r2scan | 368,502 | Inorganic crystal | r2SCAN | 40 | [23] |
| MP-ALOE | matpes_r2scan | 864,303 | Inorganic crystal | r2SCAN | 15 | [31] |
@@
The OC20 database consists of relaxation trajectories, rattled structures, and ab initio molecular dynamics (MD) configurations. For the relaxation trajectories, we employ the OC20M split provided in OC20, while for the rattled and MD structures, we use subsampled datasets. See Methods section for detailed subsampling criteria.
@@ 2.2 Frontier uMLIPs
Throughout this paper, we compare the performance of 7net-Omni with other top-performing multi-task uMLIPs, namely UMA (UMA-m-1p1) [8] and DPA (DPA-3.1-OpenLAM) [9], as well as single-task uMLIPs: eSEN (eSEN-30M-OAM, eSEN-30M-OMAT) [12], ORB (orb-v3-conservative-inf-mpa, orb-v3-conservative-inf-omat) [13], 7net-ompa (SevenNet-MF-ompa), NequIP (Nequip-OAM-L) [14], GRACE (GRACE-2L-OAM-L, GRACE-2L-OMAT-L) [15], and MACE (MACE-mpa-0-medium, MACE-omat-0-medium) [54, 55]. The specific model names tested in this work are indicated in parentheses and these correspond to the most accurate variants within each model family, as far as we can determine. We note that all the models are equivariant graph neural networks except for DPA-3.1 which utilizes invariant features. ORB employs equivariant vector features but does not strictly enforce rotational equivariance throughout the model. All the models are conservative, producing atomic forces as gradients of the energies.
We classify UMA and DPA as multi-task uMLIPs, as they are trained concurrently on databases spanning diverse material classes, including inorganic bulk, surfaces, and organic molecules. UMA is trained on OMat24, OC20, ODAC25 [32], OMC25 [33], and OMol25, whereas DPA is trained on OpenLAM-v1, which aggregates 31 databases, including large-scale datasets such as OMat24, MPtrj, OC20, and SPICE, along with several domain-specific databases. Architecturally, UMA incorporates additional embedding layers that encode information about the DFT task, charge, and spin, which are appended to the node features at each graph convolution layer. By contrast, DPA employs a one-hot encoding scheme at the final multilayer perceptron stage to distinguish training databases when predicting atomic energies. In terms of regularization, UMA uses the AdamW optimizer with a weight decay of applied to all model parameters, while DPA adopts the Adam optimizer without regularization. Thus, neither UMA nor DPA introduces the task-specific regularization proposed in this work.
Single-task uMLIP models, including eSEN, ORB, NequIP, GRACE, and MACE, typically have two variants depending on the choice of database: OMat24 or MPtrj/Alexandria, which were generated using slightly different pseudopotentials [22]. In most cases, the models are first pretrained on the OMat24 database to capture broad, high-energy configurations (e.g., eSEN-30M-OMAT). These OMat24-pretrained uMLIPs are then fine-tuned on the MPtrj and Alexandria databases (e.g., eSEN-30M-OAM) to ensure compatibility with widely used benchmarks such as Matbench Discovery [43].

### v2

@@ Abstract
Accurate yet transferable machine-learning interatomic potentials (MLIPs) are essential for accelerating materials and chemical discovery. However, many universal MLIPs overfit to chemically narrow datasets or computational protocols, limiting their reliability across diverse chemical and functional domains. We introduce a transferable multi-domain training strategy that jointly optimizes universal and task-specific parameters through selective regularization, coupled with a domain-bridging set (DBS) that aligns potential-energy surfaces across datasets. Systematic ablation experiments show that small DBS fractions (0.1%) and targeted regularization synergistically enhance out-of-distribution generalization while preserving in-domain fidelity. Trained on 15 open datasets spanning molecules, crystals, and surfaces, our model, SevenNet-Omni, achieves state-of-the-art accuracy in cross-domain scenarios, including adsorption-energy errors below 0.06 eV on metallic surfaces and 0.1 eV on metal-organic frameworks. Despite containing only 0.5% r2SCAN data, SevenNet-Omni reproduces high-fidelity r2SCAN energetics, demonstrating effective cross-functional transfer from large PBE datasets. This framework offers a scalable route toward more universal, transferable MLIPs that bridge quantum-mechanical fidelities and chemical domains.
@@ 1 Introduction
By enabling large-scale atomistic simulations with accuracy approaching that of ab initio methods such as density functional theory (DFT), machine-learning interatomic potentials (MLIPs) have significantly expanded the scope of computational materials science [1, 2, 3, 4, 5]. Recently, pretrained universal MLIPs (uMLIPs) have attracted considerable attention because they can bypass the time-consuming creation of ab initio datasets for model training [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The large databases and deep neural-network architectures behind these models enable reasonable accuracy outside the training domain [16, 17]. Currently, a multitude of large-scale ab initio databases are publicly available for training uMLIPs, covering most material classes [7, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Representative databases are shown in Fig. 1a, which indicates that the databases usually focus on specific chemical domains, reflecting distinct community interests. Each database is obtained with a unique set of computational protocols such as exchange-correlation (XC) functional, program, and ionic potentials.
@@
Many current uMLIPs were trained on datasets from specific chemical domains (e.g., inorganic crystals [16, 13, 15, 14] or organic molecules [33, 26]), limiting their accuracy in untrained material spaces. As materials engineering advances, the demand for realistic simulations that span multiple chemical domains is rapidly increasing. Examples include catalytic reactions on metal surfaces in aqueous environments [34, 35]; atomic layer deposition (ALD) in semiconductor processing [36]; the formation and evolution of the solid-electrolyte interphase [37, 38, 39]. These multi-domain simulations require uMLIPs that can deliver consistent accuracy across chemical domains by being trained over corresponding databases. However, this poses significant challenges in model training because each database is constructed with different computational protocols, which lead to distinct potential energy surfaces (PESs).
One strategy to incorporate heterogeneous datasets is to align the energy references by shifting and scaling the energies in each database separately [40, 41]. However, simply combining heterogeneous databases, even with element-dependent linear transformations, can introduce substantial noise. For example, in Supplementary Fig. 1, the PES of molecular rotation and the water dimer, calculated using various XC functionals, exhibit markedly different profiles. These non-linear discrepancies imply that PES alignment cannot be achieved solely through linear transformations. Hence, a multi-task training framework is required, in which each database is treated as a distinct task to preserve its intrinsic PES characteristics while enabling shared representation learning across domains. So far only a few uMLIPs such as DPA-3.1 [9] and UMA [8], are trained with the diverse databases concurrently within the multi-task framework.
@@
Because uMLIPs must reproduce energies consistent with a chosen set of computational parameters (e.g., XC functional or basis set), transfer learning to bridge differences across computational settings and chemical domains is essential for training multi-task uMLIPs. However, most prior work, including widely used benchmarks, has evaluated multi-task uMLIPs within single-domain scenarios rather than in complex multi-domain applications [42, 9, 8]; consequently, efficient methods for maximizing cross-domain knowledge transfer remain underexplored.
@@ 2.1 Training strategies for cross-domain generalization
In this study, we address the training of MLIPs using multiple ab initio databases that differ both in the types of atomic configurations (e.g., inorganic versus organic systems) and in the underlying computational approaches (e.g., semilocal versus hybrid functionals). To construct a single MLIP capable of accommodating databases generated under heterogeneous computational settings, we employ a multi-task MLIP framework, where each task corresponds to a different database. (In our original paper, we referred to this approach as multi-fidelity [43]. Since the present datasets often share the same level of fidelity, such as Perdew-Burke-Ernzerhof (PBE) [44] and revised PBE (RPBE) [45] within the generalized gradient approximation (GGA), we now use the term multi-task for the same methodological approach.) Within this framework, the model parameters are divided into two categories: (i) shared parameters () that are universally applied across all databases, and (ii) task-specific parameters () that are optimized exclusively for the database corresponding to task . Formally,
@@
collects higher-order terms. The right-hand side of Eq. 2 thus separates the contributions that do and do not depend on . Namely, depends solely on the shared parameters and is therefore referred to as the common PES. The second term represents the task-specific contribution.
@@
To address this issue, we introduce regularization on . By penalizing large values of , regularization encourages the model to rely more heavily on the common PES, ensuring that the shared representation captures the essential bonding characteristics across both databases. This is illustrated schematically in the central panel of Fig. 1c, where the common PES now remains close to the ground-truth PESs of both tasks, thereby enhancing knowledge transfer between the two databases.
While the regularization of is effective in improving transfer learning, it tends to fit the PES to the ground truth of in the domain. This limitation arises from the complete absence of at . To address this, we introduce a domain-bridging set (DBS), which augments by evaluating a subset of with the ab initio method used for (see the right panel of Fig. 1c). Incorporating a DBS requires only marginal computational effort (see below) yet yields substantial improvements in the accuracy of task at .
To demonstrate the effectiveness of the proposed approach, we employ the MPtrj [7] and SPICE [27] databases for MLIP training. The MPtrj database, representative of inorganic bulk systems, is generated using the PBE functional, whereas the SPICE database, curated for organic molecules, is based on hybrid functional of B97M [46]. For multi-task modeling, we adopt the SevenNet-MF architecture [43] and evaluate model performance by computing the force mean absolute error (MAE) relative to ab initio reference results.
The evaluation results are presented in Fig. 1d. As a baseline, we trained a single-task model using only the MPtrj database (left region), which exhibits large errors for molecular structures evaluated at the PBE level (denoted as Molecule@GGA). We then examined the impact of different regularization strategies in SevenNet-MF training, comparing a multi-task MLIP without regularization, a model trained with regularization applied to both and , and a model trained with regularization applied only to . As shown in the middle region of Fig. 1d, all models achieve comparable force accuracy within their respective training domains and computational settings (Crystal@GGA and Molecule@Hybrid), regardless of the regularization scheme. However, the force MAE differs significantly when evaluated on out-of-domain organic molecular configurations at the PBE level (Molecule@GGA), while all multi-task MLIPs outperform the single-task baseline. In particular, selective regularization of yields greater improvements than conventional regularization applied to all model parameters, underscoring its advantage for optimizing cross-database transferability. The right axis shows that model accuracy systematically improves as the L2 norm of decreases, indicating that effective knowledge transfer is facilitated by suppressing task-specific contributions.
@@
Building on these observations, we developed a uMLIP, 7net-Omni, by concurrently training the SevenNet-MF architecture on 15 publicly available databases (covering 13 different computational protocols) with selective task regularization and DBS. The training set spans a diverse collection of ab initio databases encompassing molecular, crystalline, metal-organic framework (MOF), and surface systems, as schematically illustrated in Fig. 1a. The specific databases, along with their chemical domains and XC functionals, are listed in Table 1. The present uMLIP concerns the charge neural system so charged structures in SPICE, OMol25 and QCML are excluded. Each database is assigned to a distinct task, but databases with identical computational protocols are grouped under the same task. We also use the term channel to refer to task-specific inference.
To construct DBS, we adopted the computational settings used in the MPtrj database and performed single-point calculations on approximately 0.1% of structures from six representative databases (see Methods for details of the sampling criterion). The numbers of sampled structures from each database are summarized in Supplementary Table 1.
To efficiently train on heterogeneous datasets, we adopted a curriculum learning strategy. Specifically, the model was first trained on crystal databases such as MPtrj, sAlex, and OMat24 to establish foundational chemical knowledge. The training was then extended to molecular systems by introducing OMol25 into the dataset. Finally, the model was trained on the whole database listed in Table 1. This curriculum-based training process enables the model to effectively capture complex chemical environments, such as adsorbate-slab systems, by utilizing prior knowledge established for crystalline and molecular systems. When a new database is introduced as an additional task, the corresponding PES is adapted by initializing the energy shift parameters via linear regression and setting them as trainable. This approach allows for a stable and coherent learning process. Further details of dataset composition, sampling ratios, and training procedures are provided in the Methods section.
@@
To examine whether the model encodes computational methods in an explainable way, we analyzed the trained parameters of 7net-Omni associated with task embeddings. The latent vectors were visualized using principal component analysis (PCA) in Fig. 1b. We find that the task-dependent weights form distinct clusters that primarily correlate with the choice of XC functional and the application of a Hubbard correction [47]. For example, molecular databases employing hybrid functionals group together in a similar region. This indicates that the model captures both similarities and discrepancies among different levels of theory, thereby contributing to the enhanced generalizability of the multi-task framework. Comparable analyses have been reported in the context of multi-task property prediction using the multiXC-QM9 dataset [48, 49]. However, unlike the present observation, it produced mixed results; for instance, embeddings generated with the same XC functional but different basis sets exhibited low similarity. This difference may stem from the use of selective task regularization in 7net-Omni, which allows the model to better recognize cross-theory similarities.
@@
| Database abbreviation | Task (Channel) | Number of structures/103 | Domain | XC functional | Over- sampling | Reference |
| MPtrj | mpa | 1,580 | Inorganic crystal | PBE(+) | 15 | [7] |
| Alex1 | mpa | 12,069 | Inorganic crystal | PBE(+) | 4 | [18, 19, 20] |
| DBS | mpa | 121 | General | PBE(+) | 15 | - |
| OMat24 | omat24 | 101,901 | Inorganic crystal | PBE(+) | 1 | [21] |
| MatPES | matpes | 412 | Inorganic crystal | PBE | 40 | [22] |
| OC202 | oc20 | 30,757 | Catalyst (metal) | RPBE | 2 | [23] |
| OC22 | oc22 | 8,210 | Catalyst (oxide) | PBE(+) | 6 | [24] |
| ODAC23 | odac23 | 4,082 | MOF | PBE-D3 | 1 | [25] |
| OMol25 (low)3 | omol25 | 60,852 | Molecule | B97M-V | 1 (5)3 | [26] |
| OMol25 (high)3 | omol25_high | 1,390 | Molecule | B97M-V | 5 | [26] |
| SPICE4 | spice | 1,738 | Molecule | B97M | 5 | [27] |
| QCML | qcml | 18,301 | Molecule | PBE0 [50] | 1 | [28] |
| MAD | mad | 86 | General | PBEsol [51] | 100 | [29] |
| MP-r2SCAN | mp_r2scan | 50 | Inorganic crystal | r2SCAN | 40 | [52, 53] |
| MatPES-r2SCAN | matpes_r2scan | 368 | Inorganic crystal | r2SCAN | 40 | [22] |
| MP-ALOE | matpes_r2scan | 864 | Inorganic crystal | r2SCAN | 15 | [30] |
@@
The OC20 database consists of relaxation trajectories, rattled structures, and ab initio molecular dynamics configurations. For the relaxation trajectories, we employ the OC20M split provided in OC20, while for the rattled and MD structures, we use subsampled datasets. See Methods section for detailed subsampling criteria.
@@ 2.2 Frontier uMLIPs
Throughout this paper, we compare the performance of 7net-Omni with other top-performing multi-task uMLIPs, namely UMA (UMA-m-1p1) [8] and DPA (DPA-3.1-OpenLAM) [9], as well as single-task uMLIPs: eSEN (eSEN-30M-OAM, eSEN-30M-OMAT) [12], ORB (orb-v3-conservative-inf-mpa, orb-v3-conservative-inf-omat) [13], 7net-ompa (SevenNet-MF-ompa), NequIP (Nequip-OAM-L) [54, 14], GRACE (GRACE-2L-OAM-L, GRACE-2L-OMAT-L) [15], and MACE (MACE-mpa-0-medium, MACE-omat-0-medium) [55, 56]. The specific model names tested in this work are indicated in parentheses and these correspond to the most accurate variants within each model family, as far as we can determine. We note that all the models are equivariant graph neural networks except for DPA-3.1 which utilizes invariant features. ORB employs equivariant vector features but does not strictly enforce rotational equivariance throughout the model. All the models are conservative, producing atomic forces as gradients of the energies.
We classify UMA and DPA as multi-task uMLIPs, as they are trained concurrently on databases spanning diverse material classes, including inorganic bulk, surfaces, and organic molecules. UMA is trained on OMat24, OC20, ODAC25 [31], OMC25 [32], and OMol25, whereas DPA is trained on OpenLAM-v1, which aggregates 31 databases, including large-scale datasets such as OMat24, MPtrj, OC20, and SPICE, along with several domain-specific databases. Architecturally, UMA incorporates additional embedding layers that encode information about the DFT task, charge, and spin, which are appended to the node features at each graph convolution layer. By contrast, DPA employs a one-hot encoding scheme at the final multilayer perceptron stage to distinguish training databases when predicting atomic energies. In terms of regularization, UMA uses the AdamW optimizer with a small weight decay applied to all model parameters, while DPA adopts the Adam optimizer without regularization. Thus, neither UMA nor DPA introduced the task-specific regularization proposed in this work.
Single-task uMLIP models, including eSEN, ORB, NequIP, GRACE, and MACE, typically have two variants depending on the choice of database: OMat24 or MPtrj/Alexandria, which were generated using slightly different pseudopotential sets [21]. In most cases, the models are first pretrained on the OMat24 database to capture broad, high-energy configurations (e.g., eSEN-30M-OMAT). These OMat24-pretrained uMLIPs are then fine-tuned on the MPtrj and Alexandria databases (e.g., eSEN-30M-OAM).